# The Role of Deep Learning Regularizations on Actors in Offline RL

### v2

@@ Abstract
Deep learning regularization techniques, such as dropout, layer normalization, or weight decay, are widely adopted in the construction of modern artificial neural networks, often resulting in more robust training processes and improved generalization capabilities. However, in the domain of Reinforcement Learning (RL), the application of these techniques has been limited, usually applied to value function estimators (Hiraoka et al., 2021; Smith et al., 2022), and may result in detrimental effects. This issue is even more pronounced in offline RL settings, which bear greater similarity to supervised learning but have received less attention. Recent work in continuous offline RL has demonstrated that while we can build sufficiently powerful critic networks, the generalization of actor networks remains a bottleneck. In this study, we empirically show that applying standard regularization techniques to actor networks in offline RL actor-critic algorithms yields improvements of 6% on average across two algorithms and three different continuous D4RL domains. [^1] Our code is available at https://github.com/DT6A/ActoReg
@@ 1 Introduction
Motivated by the above considerations, this work aims to evaluate whether DL regularization techniques offer benefits when applied to actor networks in the continuous deep offline RL setting. Drawing inspiration from Park et al. (2024) and Tarasov et al. (2024b), we select the D4RL benchmark (Fu et al., 2020) and utilize the ReBRAC (Tarasov et al., 2023) and IQL (Kostrikov et al., 2021) algorithms for our experiments.
The questions that we seek answers to in this work are as follows:
  • Are DL regularization techniques beneficial for offline RL? Answering this question, we reduce the existing gap in the application of various regularizations to RL.
  • How can different regularization techniques be effectively combined? It is known that in RL, a combination of multiple tricks might lead to further performance boost (Hessel et al., 2018, 2021; Tarasov et al., 2023).
@@ 2.2 Deep Learning regularization techniques
  • GroupNorm (Wu and He, 2018): Divides channels into groups and normalizes each group independently, providing a balance between instance normalization and batch normalization.
@@
We chose LayerNorm, FeatureNorm and GroupNorm due to their implementation simplicity and the fact that they are all similar to LayerNorm which was shown to be a very useful modification in offline RL (Tarasov et al., 2023). The choice of SpectralNorm is motivated by its successful application in another RL domain (Gogianu et al., 2021).
@@
Objective Noise (BCN) (Wang and Principe, 1999; Müller et al., 2019). Objective noise is a regularization technique where noise is added to the target values used in the loss function during training. This approach involves perturbing the target outputs slightly before computing the loss, which can help in preventing the model from overfitting to specific target values and improve its generalization capabilities.
@@ 3.1 General experimental design
Algorithms. To enable a study with sufficient evidence, we select two competitive, ensemble-free algorithms: ReBRAC (Tarasov et al., 2023), a minimalist policy-regularized approach, and IQL (Kostrikov et al., 2021), a powerful algorithm from the implicit regularization family. We use the ReBRAC + CE + MT modification of Tarasov et al. (2024a), where ReBRAC is augmented with a cross-entropy objective (Farebrother et al., 2024), leading to significant performance improvements and improved training stability on various tasks. For brevity, this modification is simply referred to as ReBRAC until specified. The choice of these two algorithms is motivated by the findings in Park et al. (2024) and Tarasov et al. (2024b), where the regular policy methods, such as ReBRAC, were identified as the most promising for offline RL, and where IQL was recognized as a strong ensemble-free alternative. When modifying ReBRAC, we implement large batch optimization (Nikulin et al., 2022) for Adroit tasks, ensuring that this aspect of ReBRAC is consistent in all tasks. It is important to note that the original ReBRAC and it’s ReBRAC+CE upgrade did not maintain this level of hyperparameter consistency; in subsection 4.3 and subsection 4.4 we compare modifications against ReBRAC+CE with large batch on Adroit. Additionally, for the IQL baseline, we remove dropout for Adroit tasks to better isolate and examine the impact of the regularization techniques. Due to computational constraints and ReBRAC’s performance, we primarily focus our experiments on it in subsection 4.2.
@@ 3.2 Regularizations leveraging
Dropout and Normalizations. In our experiments we place dropout and normalization layers after the activation function of each hidden layer within the actor network. When both regularizations are applied simultaneously, dropout is applied after the normalization, which is a conventional approach.
@@ 4.1 Which regularizations independently help actors?
In Appendix C, we present these metrics in relation to the choice of hyperparameters for each regularization technique, allowing us to draw several key conclusions. LayerNorm proves to be the most effective Layer Normalization technique; it outperforms other normalization methods in IQL and matches their performance in ReBRAC. ReBRAC is relatively insensitive to the choice of the weight decay hyperparameter $\omega$, but its performance degrades significantly when the penalty is too large. In contrast, IQL benefits from higher values of $\omega$.
@@ 4.2 Can we combine regularizations?
In this set of experiments, we explore whether combining multiple regularization techniques can lead to better performance compared to using them individually. We present the RAR rliable metrics, including the median, interquartile mean (IQM), mean, and optimality gap scores in Figure 2, with probabilities of improvement detailed in Appendix D. We use the following notation for the modifications: Algorithm+MODs+TUNED, where MODs are modifications with already tuned hyperparameters when they were applied without TUNED and TUNED is the addition to the modifications prefix who’s parameters are tuned. This experimental setup allows us to identify whether regularizations run into conflicts with each other and require additional hyperparameters tuning when combined.
@@ 4.3 Per-dataset tuning
Figure 3: rliable metrics for last checkpoint performance averaged over all Gym-MuJoCo, AntMaze and Adroit datasets when regularizations hyperparameters are tuned per-dataset. Top row Median, IQM, Mean and Optimality Gap, bottom Probability of Improvement and Performance Profiles. Every task evaluated over 10 seeds.
@@
Following the evaluation protocol outlined by Tarasov et al. (2023), we tuned hyperparameters using a different set of random seeds than those used for final evaluation. This approach ensures a fair comparison by preventing overfitting to specific random seeds or initializations.
We report the final checkpoint performance in Figure 3 and Table 1. The results clearly demonstrate that regularizations consistently benefit algorithm performance on sparse tasks (AntMaze) and tasks with a very high-dimensional action spaces and limited state-action space coverage (Adroit). Adroit results demonstrate that regularizations may help to reduce the overfitting problem when it occurs in offline RL. At the same time, the performance remains the same on the simpler tasks (Gym-MuJoCo). While combined regularizations tend to yield more substantial improvements, in some cases, a single modification may offer greater gains. Notably, we achieved a state-of-the-art score of 92.7 on AntMaze tasks, outperforming all known offline RL approaches, using the enhanced minimalist approach ReBRAC. Per-dataset results and comparisons with other algorithms are detailed in Appendix E.
@@ 4.4 Actor generalization
Next, we aimed to determine whether regularizations enhance the generalization abilities of the algorithms. To assess this, we introduced Gaussian noise during inference to either the actions produced by the actor ($\sigma=0.2$) or the observations ($\sigma=0.05$). Adding noise to the actions simulates the impact of deploying the policy in a more stochastic environment, while noise in the states evaluates the actor’s robustness to input perturbations. The results of these experiments are presented in Figure 4, with per-domain results available in Appendix F.
@@ 4.5 What is happening inside actor?
In this section, we explore the internal changes within the actor network when various regularization techniques are applied. We monitor several metrics by comparing them between training and validation sets, focusing on the output of the penultimate linear activation layer. These differences are of particular interest as they reveal how the actor perceives potential out-of-distribution data. The results are illustrated in Figure 5. We also track the network plasticity by optimizing Behaviour Clonning term using the validation set. Per-domain results are shown in Appendix G.
@@ 5 Related Work
Offline Reinforcement Learning (RL) (Levine et al., 2020) is a rapidly advancing field with numerous approaches being introduced to address the challenges of out-of-distribution state-action pairs and policy learning from static datasets. Prominent methods such as CQL (Kumar et al., 2020b), IQL (Kostrikov et al., 2021), TD3+BC (Fujimoto and Gu, 2021), and others (An et al., 2021; Chen et al., 2021; Akimov et al., 2022; Yang et al., 2022; Ghasemipour et al., 2022; Nikulin et al., 2023) have progressively tackled these challenges with increasingly sophisticated solutions. However, as these methods have evolved, they have also become more nuanced and complex.
ReBRAC (Tarasov et al., 2023) demonstrated that augmenting a minimalist offline RL approach TD3+BC (Fujimoto and Gu, 2021) with straightforward design choices can yield state-of-the-art performance across a variety of tasks. Notably, most of ReBRAC’s enhancements focused on the critic component of the actor-critic architecture. Inspired by this work and the findings of Park et al. (2024), which highlighted the actor network as a potential bottleneck in offline RL, our research explores the impact of applying popular DL regularization techniques specifically to actor networks, aiming to unlock their potential benefits.
DL regularization techniques have been extensively studied in the context of supervised learning (Kukačka et al., 2017; Smith, 2018; Tian and Zhang, 2022), where their use has become standard practice in the development of robust models. However, their application to RL has been limited. For instance, Hiraoka et al. (2021) explored the application of dropout and LayerNorm in RL, showing performance improvements, and Smith et al. (2022) found that such modifications were critical in real-world tasks. Additionally, Bhatt et al. (2019) revealed that the application of BatchNorm in RL is nuanced, with performance gains not easily achievable through straightforward implementation. In the context of offline RL, a few studies have shown that applying layer normalizations to the critic network can be beneficial (Kumar et al., 2022; Nikulin et al., 2022; Ball et al., 2023; Tarasov et al., 2023). However, all mentioned works focus on critic modifications and largely restrict their exploration to dropout and LayerNorm. To our knowledge, IQL (Kostrikov et al., 2021) is the only notable method where dropout is applied to the actor network in some tasks, yet no analysis of this application has been provided.
@@ Appendix A Experimental details
For results in subsection 4.3, we conducted a hyperparameter search and selected the best results using RAR for each dataset. We used JAX implementations of ReBRAC+CE and IQL from Tarasov et al. [2024a] based on the Clean Offline RL [Tarasov et al., 2024b] ReBRAC implementation.
@@
Our study utilized the v2 version of datasets for Gym-MuJoCo and AntMaze, and v1 for Adroit. The agents were trained for one million steps in all domains and evaluated over ten episodes for Gym-MuJoCo and Adroit and over one hundred episodes for AntMaze. Following Chen et al. [2022], AntMaze reward function is multiplied by 100.
@@
When measuring the actor plasticity loss for ReBRAC and IQL we use validation data with Behaviour Cloning (BC) term running over 100 update steps. The final loss value is a metric that we use in subsection 4.5.
Appendix B Hyperparameters
B.1 ReBRAC+CE

### v3

@@ Abstract
Deep learning regularization techniques, such as dropout, layer normalization, or weight decay, are widely adopted in the construction of modern artificial neural networks, often resulting in more robust training processes and improved generalization capabilities. However, in the domain of Reinforcement Learning (RL), the application of these techniques has been limited, usually applied to value function estimators (Hiraoka et al., 2021; Smith et al., 2022), and may result in detrimental effects. This issue is even more pronounced in offline RL settings, which bear greater similarity to supervised learning but have received less attention. Recent work in continuous offline RL (Park et al., 2024) has demonstrated that while we can build sufficiently powerful critic networks, the generalization of actor networks remains a bottleneck. In this study, we empirically show that applying standard regularization techniques to actor networks in offline RL actor-critic algorithms yields improvements of 6% on average across two algorithms and three different continuous D4RL domains. [^1] Our code is available at https://github.com/DT6A/ActoReg
@@ 1 Introduction
Motivated by the above considerations, this work aims to evaluate whether DL regularization techniques offer benefits when applied to actor networks in the continuous deep offline RL setting. Drawing inspiration from Park et al. (2024) and Tarasov et al. (2024c), we select the D4RL benchmark (Fu et al., 2020) and utilize the ReBRAC (Tarasov et al., 2024b) and IQL (Kostrikov et al., 2021) algorithms for our experiments.
In this work, we aim to answer the following questions:
  • Are DL regularization techniques beneficial for offline RL? By answering this question, we reduce the existing gap in the application of various regularizations to RL.
  • How can different regularization techniques be effectively combined? It is known that in RL, a combination of multiple tricks might lead to further performance boost (Hessel et al., 2018; 2021; Tarasov et al., 2024b).
@@ 2.2 Deep Learning regularization techniques
  • GroupNorm (Wu & He, 2018): Divides channels into groups and normalizes each group independently, providing a balance between instance normalization and batch normalization.
@@
We chose LayerNorm, FeatureNorm and GroupNorm due to their implementation simplicity and the fact that they are all similar to LayerNorm which was shown to be a very useful modification in offline RL (Tarasov et al., 2024b). The choice of SpectralNorm is motivated by its successful application in another RL domain (Gogianu et al., 2021).
@@
Objective Noise (BCN) (Wang & Principe, 1999; Müller et al., 2019). Objective noise is a regularization technique where noise is added to the target values used in the loss function during training. This approach involves perturbing the target outputs slightly before computing the loss, which can help in preventing the model from overfitting to specific target values and improve its generalization capabilities.
@@ 3.1 General experimental design
Algorithms. To enable a study with sufficient evidence, we select two competitive, ensemble-free algorithms: ReBRAC (Tarasov et al., 2024b), a modification of minimalist TD3+BC (Fujimoto & Gu, 2021) policy-regularized approach, and IQL (Kostrikov et al., 2021), a powerful algorithm from the implicit regularization family. We use the ReBRAC + CE + MT modification of Tarasov et al. (2024a), where ReBRAC is augmented with a cross-entropy objective (Farebrother et al., 2024), leading to significant performance improvements and improved training stability on various tasks. For brevity, this modification is simply referred to as ReBRAC until specified. The choice of these two algorithms is motivated by the findings in Park et al. (2024) and Tarasov et al. (2024c), where the policy-regularized methods, such as ReBRAC, were identified as the most promising for offline RL, and where IQL was recognized as a strong ensemble-free alternative. When modifying ReBRAC, we implement large batch optimization (Nikulin et al., 2022) for Adroit tasks, ensuring that this aspect of ReBRAC is consistent in all tasks. It is important to note that the original ReBRAC and it’s ReBRAC+CE upgrade did not maintain this level of hyperparameter consistency; in subsection 4.3 and subsection 4.4 we compare modifications against ReBRAC+CE with large batch on Adroit. Additionally, for the IQL baseline, we remove dropout for Adroit tasks to better isolate and examine the impact of the regularization techniques. Due to computational constraints and ReBRAC’s performance, we primarily focus our experiments on it in subsection 4.2. More details on algorithms are provided in Appendix B.
@@ 3.2 Regularizations leveraging
We provide pseudocodes for the regularizations usages in Appendix C.
Dropout and Normalizations. In our experiments we place dropout and normalization layers after the activation function of each hidden layer within the actor network. When both regularizations are applied simultaneously, dropout is applied after the normalization, which is a conventional approach. Note, that we do not combine normalizations with each other and apply only one if any.
@@ 4.1 Which regularizations independently help actors?
In Appendix E, we present these metrics in relation to the choice of hyperparameters for each regularization technique, allowing us to draw several key conclusions. LayerNorm proves to be the most effective Layer Normalization technique; it outperforms other normalization methods in IQL and matches their performance in ReBRAC. ReBRAC is relatively insensitive to the choice of the weight decay hyperparameter $\omega$, but its performance degrades significantly when the penalty is too large. In contrast, IQL benefits from higher values of $\omega$.
@@ 4.2 Can we combine regularizations?
In this set of experiments, we explore whether combining multiple regularization techniques can lead to better performance compared to using them individually. We present the RAR rliable metrics, including the median, interquartile mean (IQM), mean, and optimality gap scores in Figure 2, with probabilities of improvement detailed in Appendix F. We use the following notation for the modifications: Algorithm+MODs+TUNED, where MODs are modifications with already tuned hyperparameters when they were applied without TUNED and TUNED is the addition to the modifications prefix who’s parameters are tuned. This experimental setup allows us to identify whether regularizations run into conflicts with each other and require additional hyperparameters tuning when combined.
@@ 4.3 Per-dataset tuning
Figure 3: rliable metrics for last checkpoint performance averaged over all Gym-MuJoCo, AntMaze, and Adroit datasets when regularizations hyperparameters are tuned per dataset. Top row Median, IQM, Mean and Optimality Gap, bottom Probability of Improvement, and Performance Profiles. Every task evaluated over 10 seeds.
@@
Following the evaluation protocol outlined by Tarasov et al. (2024b), we tuned hyperparameters using a different set of random seeds than those used for final evaluation. This approach ensures a fair comparison by preventing overfitting to specific random seeds or initializations.
We report the final checkpoint performance in Figure 3 and Table 1. The results clearly demonstrate that regularizations consistently benefit algorithm performance on sparse tasks (AntMaze) and tasks with a very high-dimensional action spaces and limited state-action space coverage (Adroit). Adroit results demonstrate that regularizations may help to reduce the overfitting problem when it occurs in offline RL. At the same time, the performance remains the same on the simpler tasks (Gym-MuJoCo). While combined regularizations tend to yield more substantial improvements, in some cases, a single modification may offer greater gains. Notably, we achieved a state-of-the-art score of 92.7 on AntMaze tasks, outperforming all known offline RL approaches, using the enhanced minimalist approach ReBRAC. Per-dataset results and comparisons with other algorithms are detailed in Appendix G.
@@ 4.4 Actor generalization
Next, we aimed to determine whether regularizations enhance the generalization abilities of the algorithms. To assess this, we introduced Gaussian noise during inference to either the actions produced by the actor ($\sigma=0.2$) or the observations ($\sigma=0.05$). Adding noise to the actions simulates the impact of deploying the policy in a more stochastic environment, while noise in the states evaluates the actor’s robustness to input perturbations. The results of these experiments are presented in Figure 4, with per-domain results available in Appendix H.
@@ 4.5 What is happening inside actor?
In this section, we explore the internal changes within the actor network when various regularization techniques are applied. We monitor several metrics by comparing them between training and validation sets, focusing on the output of the penultimate linear activation layer. These differences are of particular interest as they reveal how the actor perceives potential out-of-distribution data. The results are illustrated in Figure 5. We also track the network plasticity by optimizing Behaviour Clonning term using the validation set. Per-domain results are shown in Appendix I.
@@ 5 Related Work
Offline Reinforcement Learning (RL) (Levine et al., 2020) is a rapidly advancing field with numerous approaches being introduced to address the challenges of out-of-distribution state-action pairs and policy learning from static datasets. Prominent methods such as CQL (Kumar et al., 2020b), IQL (Kostrikov et al., 2021), TD3+BC (Fujimoto & Gu, 2021), and others (An et al., 2021; Chen et al., 2021; Akimov et al., 2022; Yang et al., 2022; Ghasemipour et al., 2022; Nikulin et al., 2023) have progressively tackled these challenges with increasingly sophisticated solutions. However, as these methods have evolved, they have also become more nuanced and complex.
ReBRAC (Tarasov et al., 2024b) demonstrated that augmenting a minimalist offline RL approach TD3+BC (Fujimoto & Gu, 2021) with straightforward design choices can yield state-of-the-art performance across a variety of tasks. Notably, most of ReBRAC’s enhancements focused on the critic component of the actor-critic architecture. Inspired by this work and the findings of Park et al. (2024), which highlighted the actor network as a potential bottleneck in offline RL, our research explores the impact of applying popular DL regularization techniques specifically to actor networks, aiming to unlock their potential benefits.
DL regularization techniques have been extensively studied in the context of supervised learning (Kukačka et al., 2017; Smith, 2018; Tian & Zhang, 2022), where their use has become standard practice in the development of robust models. However, their application to RL has been limited. For instance, Hiraoka et al. (2021) explored the application of dropout and LayerNorm in RL, showing performance improvements, and Smith et al. (2022) found that such modifications were critical in real-world tasks. Additionally, Bhatt et al. (2019) revealed that the application of BatchNorm in RL is nuanced, with performance gains not easily achievable through straightforward implementation. In the context of offline RL, a few studies have shown that applying layer normalizations to the critic network can be beneficial (Kumar et al., 2022; Nikulin et al., 2022; Ball et al., 2023; Tarasov et al., 2024b). However, all mentioned works focus on critic modifications and largely restrict their exploration to dropout and LayerNorm. To our knowledge, IQL (Kostrikov et al., 2021) is the only notable method where dropout is applied to the actor network in some tasks, yet no analysis of this application has been provided.
@@ Appendix A Experimental details
For results in subsection 4.3, we conducted a hyperparameter search and selected the best results using RAR for each dataset. We used JAX implementations of ReBRAC+CE and IQL from Tarasov et al. (2024a) based on the Clean Offline RL (Tarasov et al., 2024c) ReBRAC implementation. Adam Kingma & Ba (2014) optimizer is used for both algorithms
@@
Our study utilized the v2 version of datasets for Gym-MuJoCo and AntMaze, and v1 for Adroit. The agents were trained for one million steps in all domains and evaluated over ten episodes for Gym-MuJoCo and Adroit and over one hundred episodes for AntMaze. Following Chen et al. (2022), AntMaze reward function is multiplied by 100.
@@
When measuring the actor plasticity loss for ReBRAC and IQL we use validation data with Behaviour Cloning (BC) term (MSE loss) running over 100 update steps. The final loss value is a metric that we use in subsection 4.5. This approach allows us to obtain scalar values which demonstrate the ability of the trained policy to update it’s predictions using the new data.
Appendix B Algorithms overview
Here, we provide a short description of the used algorithms.
Revisited BRAC (ReBRAC). ReBRAC (Tarasov et al., 2024b) stands as a minimalist, state-of-the-art ensemble-free algorithm for both offline and offline-to-online RL. It employs MSE to penalize deviations from actions present in the dataset. Built upon TD3+BC (Fujimoto & Gu, 2021), ReBRAC incorporates several modifications that significantly enhance its performance. The Q-loss function takes the form:
$\mathbb{E}_{(s,a,s^{\prime},\hat{a^{\prime}})\sim D,a^{\prime}\sim\pi(s^{\prime})}(Q_{\theta}(s,a)-[R(s,a)+\gamma(Q_{\hat{\theta}}(s^{\prime},a^{\prime})-\beta_{1}(a^{\prime}-\hat{a^{\prime}})^{2})])^{2}$
The policy loss function is:
$\mathbb{E}_{(s,a)\sim D,a^{\prime}\sim\pi(s^{\prime})}[-Q_{\theta}(s,a^{\prime})+\beta_{2}(a^{\prime}-a)^{2})]$
$\beta_{1}$ and $\beta_{2}$ are regularization hyperparameters.
We select ReBRAC as an exemplary algorithm with policy regularization (i.e. forcing the policy to commit actions simillar to dataset) due to its high performance and simplicity.
Implicit Q-Learning (IQL). IQL (Kostrikov et al., 2021) represents another competitive offline and offline-to-online RL algorithm. Its key advantage lies in its exclusion of out-of-distribution examples during training, a departure from most other offline RL approaches. This is achieved through the training of the V-function with the following loss:
$\mathbb{E}_{(s,a)\sim D}L_{2}^{\tau}(Q_{\theta}(s,a)-V_{\phi}(s))$
where $L_{2}^{\tau}(u)=|\tau-\mathbb{I}(u<0)|u^{2}$ is an expectile loss function with parameter $\tau$.
The Q-function loss takes the form:
$\mathbb{E}_{(s,a,s^{\prime})\sim D}(Q_{\theta}(s,a)-[R(s,a)+\gamma V_{\phi}(s^{\prime})])^{2}$
And the policy loss is:
$\mathbb{E}_{(s,a)\sim D}\exp[\beta(Q_{\theta}(s,a)-V_{\phi}(s))]\log\pi(a|s)$
where $\beta$ is algorithm hyperparameter.
Regularization is achieved through a specially designed V-function loss, without explicit penalties for the policy or value functions. We opt for IQL as it is the best example of algorithms within this family.
Appendix C Pseudocodes
In this section we provide JAX-based partial code snippets used in our experiments. Some details are ommited, see our source code for more details.
Listing 1: Actors forward function with Layer Norms and Dropout in Python JAX.
1@nn.compact
2def forwrad(self, state: jax.Array):
3 layers = [
4 nn.Dense(self.hidden_dim),
5 nn.relu,
6 nn.LayerNorm() if self.ln else id,
7 nn.LayerNorm(use_bias=False, use_scale=False) if self.fn else id,
8 nn.GroupNorm() if self.gn else id,
9 nn.Dropout(rate=self.dropout),
10 ]
11 for _ in range(self.n_hiddens - 2):
12 layers += [
13 nn.Dense(self.hidden_dim),
14 nn.relu,
15 nn.LayerNorm() if self.ln else id,
16 nn.LayerNorm(use_bias=False, use_scale=False) if self.fn else id,
17 nn.GroupNorm() if self.gn else id,
18 nn.Dropout(rate=self.dropout),
19 ]
20
21 net = nn.Sequential(layers)
22
23 # Spectral norm is added only after penultimate layer if needed
24 trunk = nn.Dense(self.hidden_dim,)(net(state)) \
25 if not self.sn else (
26 nn.SpectralNorm(nn.Dense(
27 self.hidden_dim,
28 ))(net(state), update_stats=train))
29
30 last_layer = nn.Sequential(
31 [
32 nn.relu,
33 nn.LayerNorm() if self.ln else id,
34 nn.LayerNorm(use_bias=False, use_scale=False) if self.fn else id,
35 nn.GroupNorm() if self.gn else id,
36 nn.Dropout(rate=self.dropout),
37 nn.Dense(self.action_dim),
38 nn.tanh
39 ]
40 )
41 actions = last_layer(trunk)
42
43 return actions, trunk
Listing 2: General actor update in Python JAX.
1def update_actor(
2 key: jax.random.PRNGKey, actor: TrainState, critic: TrainState,
3 batch: Dict[str, jax.Array], input_noise: float, bc_noise: float, grad_noise: float,
4):
5 key, input_noise_key, bc_noise_key, grad_noise_key = jax.random.split(key, 4)
6
7 in_noise = jax.random.normal(input_noise_key, batch["states"].shape) * input_noise
8 b_noise = jax.random.normal(bc_noise_key, batch["actions"].shape) * bc_noise
9
10 def actor_loss_fn(params: jax.Array):
11 actor_outputs = actor.apply_fn(params, batch["states"] + in_noise)
12 # algorithm specific loss function with b_noise added
13 loss = ...
14 return loss
15
16 grads = jax.grad(actor_loss_fn)(actor.params)
17
18 def add_gaussian_noise(gr, noise_std, rng_key):
19 def add_noise_to_grad(g, rng_key):
20 noise = jax.random.normal(rng_key, g.shape) * noise_std / ((1 + actor.step) ** 0.55)
21 return g + noise
22
23 leaves, tree = jax.tree_util.tree_flatten(gr)
24 rng_keys = jax.random.split(rng_key, num=len(leaves))
25 rng_keys = jax.tree_util.tree_unflatten(tree, rng_keys)
26
27 noisy_grads = jax.tree_util.tree_map(lambda g, k: add_noise_to_grad(g, k), gr, rng_keys)
28 return noisy_grads
29
30 grads = add_gaussian_noise(grads, grad_noise, grad_noise_key)
31 new_actor = actor.apply_gradients(grads=grads)
32
33 return key, new_actor
Listing 3: Full code of our custom implementation of AdamW to support elastic net penalty. Note, it might be incompatible with newer JAX/Optax versions.
1AddDecayedWeightsState = base.EmptyState
2
3def add_elastic_weights(
4 weight_decay: Union[float, jax.Array] = 0.0,
5 l1_ratio: float = 0.0,
6 mask: Optional[Union[Any, Callable[[base.Params], Any]]] = None
7) -> base.GradientTransformation:
8 def init_fn(params):
9 del params
10 return AddDecayedWeightsState()
11
12 def update_fn(updates, state, params):
13 if params is None:
14 raise ValueError(base.NO_PARAMS_MSG)
15 updates = jax.tree_util.tree_map(
16 lambda g, p: g + weight_decay * ((1 - l1_ratio) * p + l1_ratio * jnp.sign(p)), updates, params)
17 return updates, state
18
19 if mask is not None:
20 return wrappers.masked(
21 base.GradientTransformation(init_fn, update_fn), mask)
22 return base.GradientTransformation(init_fn, update_fn)
23
24
25def adamw_elastic(
26 learning_rate: base.ScalarOrSchedule,
27 b1: float = 0.9,
28 b2: float = 0.999,
29 eps: float = 1e-8,
30 eps_root: float = 0.0,
31 mu_dtype: Optional[Any] = None,
32 weight_decay: float = 1e-4,
33 l1_ratio: float = 0.0,
34 mask: Optional[Union[Any, Callable[[base.Params], Any]]] = None,
35 *,
36 nesterov: bool = False,
37) -> base.GradientTransformation:
38
39 return combine.chain(
40 transform.scale_by_adam(
41 b1=b1,
42 b2=b2,
43 eps=eps,
44 eps_root=eps_root,
45 mu_dtype=mu_dtype,
46 nesterov=nesterov,
47 ),
48 add_elastic_weights(weight_decay, l1_ratio, mask),
49 transform.scale_by_learning_rate(learning_rate),
50 )
51
52# Optax usage
53optimizer = adamw_elastic(
54 learning_rate=config.lr, weight_decay=config.wd, l1_ratio=config.l1_ratio
55)
Appendix D Hyperparameters
D.1 ReBRAC+CE